# Wasserstein–Aitchison GAN for angular measures of multivariate extremes

### v1

@@ Abstract
Economically responsible mitigation of multivariate extreme risks – extreme rainfall in a large area, huge variations of many stock prices, widespread breakdowns in transportation systems – requires estimates of the probabilities that such risks will materialize in the future. This paper develops a new method, Wasserstein–Aitchison Generative Adversarial Networks (WA-GAN), which provides simulated values of future $d$-dimensional multivariate extreme events and which hence can be used to give estimates of such probabilities. The main hypothesis is that, after transforming the observations to the unit-Pareto scale, their distribution is regularly varying in the sense that the distributions of their radial and angular components (with respect to the $L_{1}$-norm) converge and become asymptotically independent as the radius gets large. The method is a combination of standard extreme value analysis modeling of the tails of the marginal distributions with nonparametric GAN modeling of the angular distribution. For the latter, the angular values are transformed to Aitchison coordinates in a full $(d-1)$-dimensional linear space, and a Wasserstein GAN is trained on these coordinates and used to generate new values. A reverse transformation is then applied to these values and gives simulated values on the original data scale. The method shows good performance compared to other existing methods in the literature, both in terms of capturing the dependence structure of the extremes in the data, as well as in generating accurate new extremes of the data distribution. The comparison is performed on simulated multivariate extremes from a logistic model in dimensions up to 50 and on a 30-dimensional financial data set.
Keywords: Extreme value theory, Wasserstein distance, Generative adversarial networks, Multivariate analysis, Aitchison coordinates
@@ 1 Introduction
Dependence between extreme events is often of crucial importance. Extreme rainfall will be even more dangerous if it occurs at the same time at many spatial locations; costs associated with breakdowns in transportation systems are more harmful if they happen in many parts of the system; and extreme fluctuations in financial markets can cause much more damage if they occur in many parts of the economy. During the last decades the statistical research on dependent, or multivariate, extremes has grown rapidly and now contains many useful methods to model dependence between extreme values.
These methods typically build on parametric statistical models. However in the last three years papers using neural networks to simulate new multivariate extremes have started to appear. Often, but not always, these papers aim at nonparametric modeling of the dependence structure of the multivariate extreme values. The end result in the papers is an algorithm which can produce simulated values from the neural network “model” for the observations. A common approach is to transform the marginal distributions to some standard scale, then to combine nonparametric neural network modeling of the dependence structure with a parametric model from extreme value analysis (EVA) to simulate values on the standard scale, and finally to apply the inverse of the marginal transform to get simulated values on the real scale. Section 2.1 below gives a brief survey of some of this literature. We refer to [3] for a more extensive theoretical and simulation based overview of the area.
@@
    Using Aitchison coordinates from compositional data, the angular values are transformed to values on the full $\mathbb{R}^{d-1}$ space.
@@
    A Wasserstein GAN is trained to use the Aitchison coordinates to produce simulated values on the $L_{1}$-sphere.
@@
The next section contains the background needed for this together with a result (Proposition 1) providing the theoretical tool to transform the angular data of step 4 above to the real GP scale as described in step 5. Section 3 describes the details of WA-GAN, and provides the algorithms which are used by it. In Section 4.2 WA-GAN is compared to HTGAN [18] and GPGAN [28] on simulated data with a logistic dependence structure in dimension $d\in\{10,50\}$ and, in Section 4.3, on financial data consisting of daily returns for $d=30$ industry portfolios. Section 5 sums up the results of this paper.
@@ 2 Background
This section starts with a brief survey of the literature on neural network modeling of multivariate extremes. Section 2.2 gives background on regular variation and EVA and develops some new theory. Sections 2.3 and 2.4 contain brief recapitulations of Wasserstein GANs and of Aitchison coordinates. These are the building blocks which we use in Section 3 to construct the WA-GAN method.
@@ 2.1 A brief literature survey
For the terminology used in this survey see Section 2.2 below and see monographs on extreme value theory such as [7], [11] and [37].
@@
In [2] a one-layer Re-LU network is combined with a new parametrization of a GAN to generate multivariate extreme data from heavy-tailed distributions. The method is called EV-GAN and is shown to be superior to a standard GAN.
@@
ExtVAE [27] is based on the assumption that the observations are multivariate regularly varying and that the latent variable in the variational autoencoder (VAE) follows an inverse Gamma distribution and, applying some further constraints to insure regular variation, simulates values of the radial component. It then uses another VAE to produce simulated angles on the unit $L_{1}$-sphere using a VAE with Dirichlet distributed latent variables. Simulated observations are then obtained by multiplying the radial variable with the the angular value.
@@
Under the assumption that estimates of the true marginal tail indices are are available, HTGAN [18] transforms multivariate observations to the unit-Pareto scale, trains a GAN with heavy-tailed latent variables and transforms simulated variables back to original scale. It generates samples from the entire data set, not just from the extreme part.
@@
GP-GAN [28] adaptively selects exceedance levels, fits univariate GP distributions to the excesses of these levels, uses a standard GAN to propose a spectral value on the exponential scale, and then uses the fitted univariate GP distributions and the spectral value to generate a multidimensional value on the original scale. This value is then sent to the discriminator for optimization of the spectral GAN.
@@
The evtGAN [8] is aimed at multivariate block maxima observations. It uses one-dimensional Generalized Extreme Value (GEV) distributions to transform marginals to approximately uniform $[0,1]$ distributions, then trains a convolutional GAN on the uniform scale, and finally applies the inverse transform of the marginal distribution to obtain simulated values on the scale of the observations. The structure of a multivariate GEV distribution is not used in the simulation of the uniform variables.
@@
The paper [21] uses the GEV model to transform observations of block maxima to the Pickands dependence function scale and then uses either the spectral measure or the Pickands measure restrictions to train a GAN to simulate from the dependence function. The simulated observations are then transformed back to original scale. The authors argue that this approach also is useful for efficient simulation from parametric models.
@@
Finally, DeepGauge [24] applies a different framework than the references above derived from the theory of geometric extremes. It uses a multi-layer perceptron with ReLu activation function to first model radial thresholds and then another multi-layer perceptron to find a gauge function which describes the boundary of the deterministic limiting shape of scaled sample clouds.
@@ 2.2 Regular variation and extreme value analysis
Let $\bm{X}=(X_{1},\ldots,X_{d})$ be the random vector of interest with values in $\mathbb{R}^{d}$. Based on a random sample of the (unknown) distribution of $\bm{X}$, we wish to generate new samples with the same tail behavior as $\bm{X}$, even at levels beyond those encountered in the sample. Such extrapolation is possible only if we are willing to make some regularity assumptions. Our core assumptions to obtain accurate modeling of the tail of $\bm{X}$ are founded in the theory of multivariate regular variation [36]. There are two approaches to use this hypothesis.
@@
Intuitively, the measure $\nu$ helps in describing the tail behavior of $\bm{V}$ as for large $t>0$, we have $\mathsf{P}(\bm{V}\in tB)\approx\nu(B)/t$ where $tB\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\{t\bm{x}:\bm{x}\in B\}$. The measure $\nu$ is often referred to as the exponent measure of $\bm{V}$, because of its appearance in the exponent of the expression of the multivariate extreme value distribution to which the law of $\bm{V}$ is attracted [11, Definition 6.1.7].
Our procedure below will benefit from an equivalent formulation of Assumption 1 in polar coordinates with respect to the $L_{1}$-norm $|\bm{x}|_{1}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sum_{j=1}^{d}|x_{j}|$ for $\bm{x}\in\mathbb{R}^{d}$. More concretely, it can be shown [37, Theorem 6.1] that due to a convenient homogeneity property of $\nu$ [11, Theorem 6.1.9], if we put
@@
$\lim_{t\rightarrow\infty}\mathsf{P}\left(\bm{W}\in A\mid R\geq t\right)=\Phi(A)$ (2)
@@
It is possible for $\Phi$ to have positive mass on one of the faces of the simplex $\Delta_{d-1}$, that is, sets for which at least one of the coordinates is zero. In terms of extremes, it corresponds to scenarios where some components of $\bm{X}$ are large while others are small. These are called extreme directions and have to be treated by specific methods which our methodology is not able to handle, see, e.g., [31]. We exclude this situation here.
@@
By the Pickands–Balkema–de Haan theorem [6, 34], Assumption 3 is equivalent to the assumption that $X_{j}$ is in the maximum domain of attraction of a GEV distribution. The limit in Assumption 3 holds uniformly in $y>0$. The assumption justifies the following approximation for $y>0$ such that $01+\xi_{j}y/\sigma_{j}>$ 0:
@@
where $\sigma_{j}>0$ is a scale parameter that accounts for the unknown function $\sigma_{j}(\cdot)$. Estimation of the bivariate parameter $(\sigma_{j},\xi_{j})$ is typically performed by the method of maximum likelihood. This is the well-known Peaks-Over-Threshold (PoT) approach for modeling the tails of real-valued data, see for instance [42] and [7, Section 5.3]. A random variable with distribution function
@@
As already explained, Assumption 3 is equivalent to the hypothesis that $X_{j}$ is in the maximum domain of attraction of a GEV distribution. As a consequence, $X_{j}$ satisfies all the equivalent formulations of this condition as presented, e.g., in Theorem 1.1.6 of [11]. As some of them will be needed below, we recall them here. Let $b_{j}(t)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}F_{j}^{-1}(1-1/t)=(1/(1-F_{j}))^{-1}(t)$, for $t>1$, be the tail quantile function of $F_{j}$. Then, the following statements are equivalent to Assumption 3.
@@
    where the right-hand side is to be interpreted as $\log y$ if $\xi_{j}=0$. The convergence also holds locally uniformly in $y\in(0,\infty)$, see Section 1.2 in [11].
@@
What do Assumptions 13 tell us about the tail behavior of $\bm{X}$? The answer can be conveniently formulated in terms of multivariate generalized Pareto (MGP) distributions, which extend Assumption 3 to the multivariate case. These distributions have been introduced in [40] and are reviewed in [32]. Below, we provide an introduction to their usage in modeling multivariate extremes along with a theoretical result (Proposition 1) which will be useful for sampling from a MGP distribution.
In arbitrary dimension, it is not clear how to define an “extreme” point. Some authors are interested in modeling the vector of componentwise maxima of the data, leading to the multivariate GEV distributions, studied extensively in the literature after the pioneering work of [12]. Recently, see e.g. [39], more attention has been given to an extension of the standard PoT procedure in a multivariate context. The approach relies on the fact that the excess of $\bm{X}$ above a large threshold vector $\bm{u}$ asymptotically follows an MGP distribution, with an appropriate definition of when an exceedance over a multivariate threshold takes place.
More concretely, given a vector $\bm{u}\in\mathbb{R}^{d}$ of “high thresholds”, i.e., below but close to $\bm{u}_{*}=(u_{*j})_{j=1}^{d}$, we consider the random vector of excesses
@@
The random vector $\bm{H}$ in (10) is MGP($\bm{\sigma},\bm{\xi},\bm{S}$) distributed, in the sense of [32, Definition 2.2], with spectral generator $\bm{S}$ having distribution
@@
where $\bm{U}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\log(\bm{\Theta})$, with $\bm{\Theta}\sim\Phi$, and $Q\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\max(\bm{U})$. In other words, the random vector $\bm{U}$ is a $U$-generator of $\bm{H}$ in the sense of Proposition 9 of [38]; see also equation (12) in [32]. Note in particular that the moment condition $0<\mathsf{E}[\exp(U_{j})]<\infty$ for all $j\in\{1,\ldots,d\}$ is satisfied as we have, from (3),
@@
Here, since we work mainly with the angular measure $\Phi$ in our procedure, we decided to consider the multivariate Pareto random vector $\bm{Y}$, as in (8), to be the “standard” MGP, that is with $\bm{\xi}=\bm{1}$ and $\bm{\sigma}=\bm{1}$ instead of the choice $\bm{\xi}=\bm{0}$ made in [38, 32, 39], for example. This approach has also been considered in [17]. Of course, it suffices to take $\bm{Z}=\log(\bm{Y})$ to obtain an MGP random vector with parameter $\bm{\xi}=\bm{0}$, as can be seen from comparing (10) to Equation (4) in [32].

### v2

@@ Abstract
Economically responsible mitigation of multivariate extreme risks—extreme rainfall in a large area, huge variations of many stock prices, widespread breakdowns in transportation systems—requires estimates of the probabilities that such risks will materialize in the future. This paper develops a new method, Wasserstein–Aitchison Generative Adversarial Networks (WA-GAN) to, which provides simulated values of $d$-dimensional multivariate extreme events and which can hence be used to give estimates of such probabilities. The main hypothesis is that, after transforming the observations to the unit-Pareto scale, their distribution is regularly varying in the sense that the distributions of their radial and angular components (with respect to the $L_{1}$-norm) converge and become asymptotically independent as the radius gets large. The method is a combination of standard extreme value analysis modeling of the tails of the marginal distributions with nonparametric GAN modeling of the angular distribution. For the latter, the angular values are transformed to Aitchison coordinates in a full (d1)(d-1)-dimensional linear space, and a Wasserstein GAN is trained on these coordinates and used to generate new values. A reverse transformation is then applied to these values and gives simulated values on the original data scale. Our method is applied to simulated data and to a financial data set from the Kenneth French Data Library. The method shows good performance compared to other existing methods in the literature, both in terms of capturing the dependence structure of the extremes in the data and in generating accurate new extremes.
Keywords: Extreme value theory, Wasserstein distance, Generative AI for extremes, Generative adversarial networks, Multivariate analysis, Aitchison coordinates
@@ 1 Introduction
The dependence between high-dimensional extreme events is often of crucial importance. Extreme rainfall will be even more dangerous if it occurs at the same time in many spatial locations; costs associated with breakdowns in transportation systems are more harmful if they occur in many parts of the system; and extreme fluctuations in financial markets can cause much more damage if they occur in many parts of the economy. During the last decades, the statistical research on dependent, or multivariate, extremes has grown rapidly and now contains many useful, mostly parametric, methods to model dependence between extreme values. Often, the computational challenge of estimating the parameters of these models restricts their use to low-dimensional situations.
However, generative artificial intelligence (AI) and neural networks now provide completely new opportunities to meet these challenges and to use much higher-dimensional extreme data than before. During the last three years, papers using these opportunities have started to appear at an increasing rate. Often, but not always, these papers aim at nonparametric modeling of the dependence structure of the multivariate extreme values. The end result in the papers is an algorithm that can produce simulated values from the neural network “model” for the observations. A common approach is to transform the marginal distributions to some standard scale, then to combine nonparametric neural network modeling of the dependence structure with a parametric model from extreme value analysis to simulate values on the standard scale, and finally to apply the inverse of the marginal transform to get simulated values on the real scale. Section 2.1 gives a brief survey of some of this literature. We refer to [Allouche-et-al] for a more extensive theoretical and simulation based overview of the area.
@@
    Using Aitchison coordinates from compositional data, the angular values are transformed into values on the whole $\mathbb{R}^{d-1}$ space.
@@
    A Wasserstein GAN with Gaussian latent variables is trained to use the Aitchison coordinates to produce simulated values on the $L_{1}$-sphere.
@@
Section 2 further contains the background needed for this together with a result (Proposition 1) providing the theoretical tool to transform the angular data on the $L_{1}$ sphere of step 4 to the real GP scale as described in step 5. Section 3 describes the details of WA-GAN, and provides the actual algorithms. In Section 4.2, WA-GAN is compared to HTGAN [Girard-et-al] and GPGAN [Li-et-al] on simulated data with a logistic and Hüsler–Reiss dependence structures in dimension $d\in\{10,20,50\}$ and, in Section 4.3, on financial data consisting of daily returns for $d=30$ industry portfolios. Section 4.2.3 describes how the use of non-Gaussian latent distribution for the generator may help improve the results when asymptotic independence is suspected. Section 5 sums up the results of this paper.
@@ 2 Background
This section starts with a brief survey of the literature on neural modeling of multivariate extremes. Section 2.2 gives background on regular variation and extreme value analysis and develops some new theory. Sections 2.3 and 2.4 contain brief recapitulations of Wasserstein GANs and of Aitchison coordinates. These are the building blocks which we use in Section 3 to construct the WA-GAN method.
@@ 2.1 A brief literature survey
For the terminology used in this survey see Section 2.2 below and the standard extreme value theory monographs [beirlant2004, deHaanFerreira2006, Resnick2007] or the more recent ones [mikosch2024evt, resnick2024hidden].
@@
In [Allouche-et-al-EV-GAN] a one-layer Re-LU network is combined with a new parametrization of a GAN to generate multivariate extreme data from heavy-tailed distributions. The method is called EV-GAN and is shown to be superior to a standard GAN. The same authors in [Allouche-et-al-excessGAN] develop an ExcessGAN method to simulate variables conditioned on all components exceeding a high level. It employs a GAN-based architecture with a specifically designed network structure aimed at reducing the bias.
@@
ExtVAE [VAE-Lafon-et-al] is based on the assumption that the observations are multivariate regularly varying and that the latent variable in the variational autoencoder (VAE) follows an inverse Gamma distribution and, applying some further constraints to insure regular variation, simulates values of the radial component. It then uses another VAE to produce simulated angles on the unit $L_{1}$-sphere using a VAE with Dirichlet distributed latent variables. Simulated observations are then obtained by multiplying the radial variable with the the angular value.
@@
Under the assumption that estimates of the true marginal tail indices are are available, HTGAN [Girard-et-al] transforms multivariate observations to the unit-Pareto scale, trains a GAN with heavy-tailed latent variables and transforms simulated variables back to original scale. It generates samples from the entire data set, not just from the extreme part.
@@
GP-GAN [Li-et-al] adaptively selects exceedance levels, fits univariate GP distributions to the excesses of these levels, uses a standard GAN to propose a spectral value on the exponential scale, and then uses the fitted univariate GP distributions and the spectral value to generate a multidimensional value on the original scale. This value is then sent to the discriminator for optimization of the spectral GAN. The authors show superior performance of their method to the ExGAN approach developed in [exGAN-bathia-et-al2021] on a benchmark US precipitation dataset.
@@
The evtGAN [Boulagiem-et-al] is aimed at multivariate block maxima observations. It uses one-dimensional Generalized Extreme Value (GEV) distributions to transform marginals to approximately uniform $[0,1]$ distributions, then trains a convolutional GAN on the uniform scale, and finally applies the inverse transform of the marginal distribution to obtain simulated values on the scale of the observations. The structure of a multivariate GEV distribution is not used in the simulation of the uniform variables.
@@
The paper [dMNNs-hasan-et-al] uses the GEV model to transform observations of block maxima to the Pickands dependence function scale and then uses either the spectral measure or the Pickands measure restrictions to train a GAN to simulate from the dependence function. The simulated observations are then transformed back to original scale. The authors argue that this approach also is useful for efficient simulation from parametric models.
@@
Finally, the methods in [hu2025gpdflowgenerativemultivariatethreshold, DeepGauge-Callum-et-al, demonte2025generative] rely on a different framework and consider the setting of “geometric extremes”. In [DeepGauge-Callum-et-al], a multi-layer perceptron with ReLu activation function is used to model radial thresholds, followed by another multi-layer perceptron to find a gauge function which describes the boundary of the deterministic limiting shape of scaled sample clouds. Alternatively, [hu2025gpdflowgenerativemultivariatethreshold] and [demonte2025generative] use normalizing flows to simulate multivariate extremes. Issues related to generating heavy-tailed outputs based on normalizing flows are also discussed in [Hikle-et-al-TFF].
A broader comparison of many existing generative methods for extremes is provided in [wessel2025comparisongenerativedeeplearning].
@@ 2.2 Regular variation and extreme value analysis
Let $\bm{X}=(X_{1},\ldots,X_{d})$ be the random vector of interest with values in $\mathbb{R}^{d}$. Based on a random sample of the (unknown) distribution of $\bm{X}$, we wish to generate new samples with the same tail behavior as $\bm{X}$, even at levels beyond those encountered in the sample. Such extrapolation is possible only if we are willing to make some regularity assumptions. Our core assumptions to obtain accurate modeling of the tail of $\bm{X}$ are founded in the theory of multivariate regular variation [resnick1987]. There are two approaches to use this hypothesis.
@@
Intuitively, the measure $\nu$ helps in describing the tail behavior of $\bm{V}$ as for large $t>0$, we have $\mathsf{P}(\bm{V}\in tB)\approx\nu(B)/t$ where $tB\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\{t\bm{x}:\bm{x}\in B\}$. The measure $\nu$ is often referred to as the exponent measure of $\bm{V}$, because of its appearance in the exponent of the expression of the multivariate extreme value distribution to which the law of $\bm{V}$ is attracted [deHaanFerreira2006, Definition 6.1.7].
Our procedure below will benefit from an equivalent formulation of Assumption 1 in polar coordinates with respect to the $L_{1}$-norm $|\bm{x}|_{1}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sum_{j=1}^{d}|x_{j}|$ for $\bm{x}\in\mathbb{R}^{d}$. More concretely, it can be shown [Resnick2007, Theorem 6.1] that due to a convenient homogeneity property of $\nu$ [deHaanFerreira2006, Theorem 6.1.9], if we put
@@
$\lim_{t\rightarrow\infty}\mathsf{P}\left(\bm{W}\in A\mid R\geq t\right)=\Phi(A),$ (2)
@@
It is possible for $\Phi$ to have positive mass on one of the faces of the simplex $\Delta_{d-1}$, that is, sets for which at least one of the coordinates is zero. In terms of extremes, it corresponds to scenarios where some components of $\bm{X}$ are large while others are small. These are called extreme directions and have to be treated by specific methods which our methodology is not able to handle, see, e.g., [mourahib2024directions]. We exclude this situation here.
@@
By the Pickands–Balkema–de Haan theorem [balkema1974residual, pickands1975statistical], Assumption 3 is equivalent to the assumption that $X_{j}$ is in the maximum domain of attraction of a GEV distribution. The limit in Assumption 3 holds uniformly in $y>0$. The assumption justifies the following approximation for $y>0$ such that $1+\xi_{j}y/\sigma_{j}>0$:
@@
where $\sigma_{j}>0$ is a scale parameter that accounts for the unknown function $\sigma_{j}(\cdot)$. Estimation of the bivariate parameter $(\sigma_{j},\xi_{j})$ is typically performed by the method of maximum likelihood. This is the well-known Peaks-Over-Threshold (PoT) approach for modeling the tails of real-valued data, see for instance [smith1987estimating] and [beirlant2004, Section 5.3]. A random variable with distribution function
@@
As already explained, Assumption 3 is equivalent to the hypothesis that $X_{j}$ is in the maximum domain of attraction of a GEV distribution. As a consequence, $X_{j}$ satisfies all the equivalent formulations of this condition as presented, e.g., in Theorem 1.1.6 of [deHaanFerreira2006]. As some of them will be needed below, we recall them here. Let $b_{j}(t)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}F_{j}^{-1}(1-1/t)=(1/(1-F_{j}))^{-1}(t)$, for $t>1$, be the tail quantile function of $F_{j}$. Then, the following statements are equivalent to Assumption 3.
@@
    where the right-hand side is to be interpreted as $\log y$ if $\xi_{j}=0$. The convergence also holds locally uniformly in $y\in(0,\infty)$, see Section 1.2 in [deHaanFerreira2006].
@@
What do Assumptions 13 tell us about the tail behavior of $\bm{X}$? The answer can be conveniently formulated in terms of multivariate generalized Pareto (MGP) distributions, which extend Assumption 3 to the multivariate case. These distributions have been introduced in [rootzen2006] and are reviewed in [Multivariate-Naveau-segers]. Below, we provide an introduction to their usage in modeling multivariate extremes along with a theoretical result (Proposition 1) which will be useful for sampling from a MGP distribution.
In arbitrary dimension, it is not clear how to define an “extreme” point. Some authors are interested in modeling the vector of componentwise maxima of the data, leading to the multivariate GEV distributions, studied extensively in the literature after the pioneering work of [dehaan1977]. Recently, see e.g. [rootzen2018], more attention has been given to an extension of the standard PoT procedure in a multivariate context. The approach relies on the fact that the excess of $\bm{X}$ above a large threshold vector $\bm{u}$ asymptotically follows an MGP distribution, with an appropriate definition of when an exceedance over a multivariate threshold takes place.
More concretely, given a vector $\bm{u}\in\mathbb{R}^{d}$ of “high thresholds”, i.e., below but close to $\bm{u}_{*}=(u_{*j})_{j=1}^{d}$, where $u_{*j}$ denotes the (possibly infinite) upper endpoint of $X_{j}$, we consider the random vector of excesses
@@
The random vector $\bm{H}$ in (10) is MGP($\bm{\sigma},\bm{\xi},\bm{S}$) distributed, in the sense of [Multivariate-Naveau-segers, Definition 2.2], with spectral generator $\bm{S}$ having distribution
@@
where $\mathbb{I}\{\mathcal{E}\}$ denotes the indicator random variable of the event $\mathcal{E}$ and $\bm{U}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\log(\bm{\Theta})$, with $\bm{\Theta}\sim\Phi$, and $Q\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\max(\bm{U})$. In other words, the random vector $\bm{U}$ is a $U$-generator of $\bm{H}$ in the sense of Proposition 9 of [rootzen2018multivariate]; see also equation (12) in [Multivariate-Naveau-segers]. Note in particular that the moment condition $0<\mathsf{E}[\exp(U_{j})]<\infty$ for all $j\in\{1,\ldots,d\}$ is satisfied as we have, from (3),
@@
Here, since we work mainly with the angular measure $\Phi$ in our procedure, we decided to consider the multivariate Pareto random vector $\bm{Y}$, as in (8), to be the “standard” MGP, that is with $\bm{\xi}=\bm{1}$ and $\bm{\sigma}=\bm{1}$ instead of the choice $\bm{\xi}=\bm{0}$ made in [rootzen2018multivariate, Multivariate-Naveau-segers, rootzen2018], for example. This approach has also been considered in [ferreira2014pareto]. Of course, it suffices to take $\bm{Z}=\log(\bm{Y})$ to obtain an MGP random vector with parameter $\bm{\xi}=\bm{0}$, as can be seen from comparing (10) to Equation (4) in [Multivariate-Naveau-segers].